OpenStudy (anonymous):

Help !! Will Fan and Medall !!:)) A mass of 37.5 g of mercury (II) oxide is decomposed by heating. a.How many Moles of mercury (II) oxide are decomposed? b.How many Grams of oxygen are prepared?

OpenStudy (anonymous):

In stoichiometry, it is important to: -Write and balance the equation. - Consider ratio of moles, masses or of volumes - The formula for moles is \[n=\frac{ m }{ Mr }\], where n is mol, m is given mass and Mr is Molecular atomic mass. Working out this problem, we firstly have \[2Hg_{2}O \rightarrow 4Hg + O_{2}\],A. they want number of moles and we have to use the aforementioned formula n=37.5/201=0.19mol B.remember to take account of ratio of moles, Hg2O to O2 are in the ratio of 4:1, meaning oxygen has a fourth of the moles of Hg2O or 0.19/4=0.0475mol Back to the formula of moles, we are to find the given mass reacted and making m the subject becomes m=n*Mr or 0.0475*(16*2)=1.52g of Oxygen.